OpenStudy (anonymous):

FInd exact value of csc ^-1 (sqrt2)

OpenStudy (anonymous):



OpenStudy (jhannybean):

I have to review this again, sorry.

OpenStudy (anonymous):

@rascal get it?

OpenStudy (jhannybean):

@RaphaelFilgueiras mind if i rewrite it using equation editor?

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

\[\alpha=arccsc(\sqrt{2})->\csc(\alpha)=\sqrt{2}->\frac{ 1 }{ \sin(\alpha) }=\sqrt{2}->\sin(\alpha)=\frac{ \sqrt{2}}{ {2} }\]

OpenStudy (anonymous):

How did you go from 1/sin(alpha)=sqrt(2) to sin(alpha)=sqrt(2)/2? I understand how you got 1/sin(alpha)=sqrt(2).

OpenStudy (anonymous):

\[->\alpha= \pi/4\]

OpenStudy (anonymous):